What is the percentage by weight of glucose if 1.52g of glucose dissolved is 6.08g of water
1
Expert's answer
2020-04-13T02:34:18-0400

ms=mgl+mH2O=1.52+6.08=7.6g

w=(mgl/ms)*100%=(1.52/7.6)*100%=16.4%


Answer: w=16.4%